# A uniquely solvable and positivity-preserving finite difference scheme for the Flory-Huggins-Cahn-Hilliard equation with dynamical boundary condition

### v1

@@ 5.1 Convergence test
Table 1 and Table 2 present the errors and convergence rates, based on the data at $T=0.001$ produced by the numerical scheme (2.15) – (2.19). The 2nd order convergence on the physical boundary is indicated in Table 1, where the inner product and the associated norm are defined in the lower dimension. It is observed that all the numerical convergence rates approach to the theoretical value of as the spatial mesh is refined.
5.2 Properties of the concentration variable
In this part, we perform some numerical simulations to verify the theoretical results, including energy decay and mass conservation. In more details, we will present several numerical examples with different initial data. In this model, we choose the parameters as: $\varepsilon=0.02,\theta_{0}=3,\kappa=1$, and the time step size and space step are taken as $h=1/128,s=10^{-5}$, respectively.
To simulate the phenomenon of spinodal decomposition, the initial data is chosen as
$\phi^{0}_{i,j}=0.2+0.02*r_{i,j},$ (5.3)
where $r_{i,j}$ are uniformly distributed random numbers in $[-1,1]$. In Figure 1 we present several snapshot evolution of $\phi$ at the selected time instants. The physical boundary is established in the left and right sides of the square domain, while a periodic boundary condition is imposed in the $y$ direction. The short-range interaction can be observed around the boundary area.
@@ 5.2 Properties of the concentration variable
In the left part of Figure 2, the evolution of energy is illustrated, which indicates an energy decay in time. We also display the evolution of mass difference of $\phi$, which is defined as $\bar{\phi}^{n}-\bar{\phi}^{0}$, with $\bar{\phi}^{n}$ given by (2.23). Such a mass difference is displayed in the right part of Figure 2, which in turn indicates the mass conservation up to a machine prevision.

### v2

@@ 5.1 Convergence test
Table 1 and Table 2 present the errors and convergence rates, based on the data at $T=0.001$ produced by the numerical scheme (2.15) – (2.19). The 2nd order convergence on the physical boundary is indicated in Table 1, where the inner product and the associated norm are defined in the lower dimension. It is observed that all the numerical convergence rates approach to the theoretical value as the spatial mesh is refined.
5.2 Spinodal decomposition
In this section, we perform a numerical simulation to verify energy decay and mass conservation. The parameters are chosen as follows: $\varepsilon=0.02,\theta_{0}=3,\kappa=1$. The time step size and space step are taken as $h=1/128,s=10^{-5}$, respectively.
To simulate the phenomenon of spinodal decomposition, the initial data in the bulk are chosen to satisfy
$\phi^{0}_{i,j}=0.2+0.02\times r_{i,j},$ (5.3)
where $r_{i,j}$ are uniformly distributed random numbers in $[-1,1]$. In Figure 1 we present several snapshot evolutions of $\phi$ at the selected time instants. The physical boundary is established on the left and right sides of the square domain, while a periodic boundary condition is imposed in the $y$ direction.
The mixture does not phase separate on the physical boundary because there are no initial fluctuations there, and $\phi$ takes on the uniform value of $0.2$ at the left and right boundaries. In the snapshots, the interface is not perpendicular to the wall, in contrast to the case where homogeneous Neumann boundary conditions are taken on the physical boundary. A transition layer at the physical boundary forms, with a certain thickness, where a short-range wetting-type interaction can be observed.
In the left part of Figure 2, the evolution of energy is illustrated, which indicates an energy decay in time. We also display the evolution of mass difference of $\phi$, which is defined as $\bar{\phi}^{n}-\bar{\phi}^{0}$, with $\bar{\phi}^{n}$ given by (2.23). The mass variation is displayed in the right part of Figure 2, and indicates mass conservation of about $1\times 10^{-9}$.
@@ 5.2 Spinodal decomposition
5.3 Droplet evolution
In this part, we initialize a square droplet with its center at point $(0.5,0.2)$, and one edge of the droplet occurs at the dynamical boundary (see Figure 3(a)). The phase value inside the droplet is $0.8$ and outside the droplet is $-0.8$, and the parameters are chosen as $\varepsilon=0.01,\theta_{0}=3,\kappa=0.01$. Figure 3 presents the evolution of the droplet. The physical, mass-conserving boundary is now at the bottom of the domain. It is observed that the square droplet gradually evolves into a circle and the interface forms a certain contact angle with the solid wall that deviates from 90 degrees. Recall that a 90-degree angle would form at a physical boundary when homogeneous Neumann boundary conditions are used. The deviation that is observed is due to the fact that mass is separately conserved on the physical boundary, and this prevents the formation of perpendicular contact angle.
(a) t=0
(b) t=0.001
(c) t=0.003
(d) t=0.005
(e) t=0.01
Figure 3: Snapshots of solution for droplet evolution at selected times.
5.4 Two droplet evolution
In this part, we numerically simulate the slow fusion process of two circular droplets. As shown in Figure 4, two droplets barely contact with each other and the bottom boundary crosses through both of them. Two numerical simulations, with homogeneous Neumann and dynamical boundary conditions, respectively, are displayed in Figure 5. In the case of homogeneous Neumann boundary condition, the interface between two kinds of liquid (yellow and blue) quickly becomes perpendicular to the boundary. Without mass conservation on the surface, the yellow part on the boundary merges into one very fast and lengthened slightly. In the case of dynamical boundary condition, the total length of the yellow part basically remains unchanged, due to the mass conservation on the surface, and the coarsening on boundary takes a longer time. The contact angle changes during the evolution process as well.
Figure 4: Initial profile of the fusion process of two circular droplets.
Figure 5: (a): The evolution of droplet fusion process, with the homogeneous Neumann boundary condition, at time instants $t=1\times 10^{-4}$, $1\times 10^{-3}$, $3\times 10^{-3}$, $1\times 10^{-2}$. (b) The evolution of droplet fusion process, with the dynamical boundary condition at the bottom boundary, at the same time instants.
5.5 Comparison of evolution rate
In section 5.4, it is observed that the dynamic boundary slows down the merging of structures near the solid physical wall. Here we perform a further comparison of evolution rate between homogeneous Neumann and dynamical boundary conditions. The initial profile is presented in Figure 6 and the simulation results are displayed in Figure 7. Once again, the physical boundary is at the bottom.
More obviously than Figure 5, the blue part surrounded by yellow gradually disappears and the mass conservation constraint greatly slows down the evolution process. We have reason to believe that, if the yellow part spans a wider range, the difference in evolutionary speed will be more pronounced. Also, at equilibrium, the contact angle deviates from 90 degrees with the mass-conserving boundary condition.
Figure 6: Initial profile of fusion process.
Figure 7: (a): The evolution for fusion process, with the homogeneous Neumann bottom boundary condition, at time instants $t=1\times 10^{-4}$, $2\times 10^{-3}$, $1\times 10^{-2}$, $3\times 10^{-2}$. (b) The evolution for fusion process, with the dynamical bottom boundary condition, at time instants $t=2\times 10^{-3}$, $3\times 10^{-2}$, $2\times 10^{-1}$, $6\times 10^{-1}$.